Integrate sin³x dx
1
Expert's answer
2020-02-18T05:33:02-0500

Solution.

"\int sin^n(x)dx = -\frac{cos(x)sin^{n-1}(x)}{n} + \frac{n-1}{n}\int sin^{-2+n}(x)dx \ , n=3"

"\int sin^3(x)dx = -\frac{1}{3}sin^2(x)cos(x) + \frac{2}{3}\int sin(x)dx = -\frac{sin^2(x)cos(x)}{3} -\frac{2cos(x)}{3} + C"

Answer:

"\frac{1}{12}(cos(3x)-9cos(x)) + C"